OpenStudy (frank0520):

Suppose a, b, c and d are non-zero constants such that a d – b c = 1. Show that the augmented matrix a b | e c d | f  always represents a consistent and independent system of linear equations, no matter what the values of e and f.

OpenStudy (frank0520):

This is what I got so far and now I'm stuck: \[\left[\begin{matrix}a & b | e \\ c & d|f\end{matrix}\right]\] \[\left[\begin{matrix}ac & bc|ec \\ ac & ad|fa\end{matrix}\right]\] \[\left[\begin{matrix}ac & bc|ec \\ 0 & ad-bc|fa-ec\end{matrix}\right]\]